Cholesterol, C27H46O, is a waxy substance found naturally in the body. Under some
circumstances cholesterol can clog the arteries. Health Canada recommends that the total
cholesterol concentration does not exceed 5.2 x 10-3 mol/L. If you have about 4.7 L of blood,
what is the maximum mass of cholesterol that your blood should contain?
1
Expert's answer
2020-05-24T02:36:52-0400

The maximum mass of cholesterol can be calculated using the following steps:

1. Calculate the number of moles of cholesterol in blood.

ncholesterol = M × V

where M - maximum cholesterol concentration, V - blood volume.

From here:

ncholesterol = 5.2 × 10-3 mol/L × 4.7 L = 2.4 × 10-2 mol

2. Calculate the molar mass of cholesterol:

Mrcholesterol = (12 g/mol × 27) + (1 g/mol × 46) + (16 g/mol × 1) = 386 g/mol

3. Calculate the mass of cholesterol:

mcholesterol = ncholesterol × Mrcholesterol = 2.4 × 10-2 mol × 386 g/mol = 9.3 g


Answer: 9.3 g